# Rigorous results on approach to thermal equilibrium, entanglement, and nonclassicality of an optical quantum field mode scattering from the elements of a non-equilibrium quantum reservoir

### v1

@@ Abstract
Rigorous derivations of the approach of individual elements of large isolated systems to a state of thermal equilibrium, starting from arbitrary initial states, are exceedingly rare. This is particularly true for quantum mechanical systems. We demonstrate here how, through a mechanism of repeated scattering, an approach to equilibrium of this type actually occurs in a specific quantum system, one that can be viewed as a natural quantum analog of several previously studied classical models. In particular, we consider an optical mode passing through a reservoir composed of a large number of sequentially-encountered modes of the same frequency, each of which it interacts with through a beam splitter. We then analyze the dependence of the asymptotic state of this mode on the assumed stationary common initial state $\sigma$ of the reservoir modes and on the transmittance $\tau=\cos\lambda$ of the beam splitters. These results allow us to establish that at small $\lambda$ such a mode will, starting from an arbitrary initial system state $\rho$, approach a state of thermal equilibrium even when the reservoir modes are not themselves initially thermalized. We show in addition that, when the initial states are pure, the asymptotic state of the optical mode is maximally entangled with the reservoir and exhibits less nonclassicality than the state of the reservoir modes.
@@ 1 Introduction
There is continued interest (see for example [1, 2, 3, 4, 5] and references therein) in the longstanding problem of how large systems, particularly quantum mechanical ones, undergo the ubiquitous process of thermalization, i.e., how it is that they are inevitably observed to approach a state of thermal equilibrium, starting from essentially arbitrary initial conditions. For large, isolated quantum mechanical systems, much of this recent interest has focused on verifying the Eigenstate Thermalization Hypothesis (ETH) of Deutsch (see [6] for an overview). According to the ETH, energy eigenstates of large systems tend, overwhelmingly, to have macroscopic properties consistent with the thermodynamically-equilibrated states that the systems are assumed to approach.
@@
In this paper we present a fully quantum mechanical model in which a similar mechanism of repeated scattering drives a single degree of freedom of a many-body system to thermal equilibrium. More precisely, we consider a single optical field mode (the “system” mode) which couples to a large reservoir of $K$ independent optical field modes of the same frequency through a sequence of $K$ identical beam splitters, each having a transmittance $\tau=\cos\lambda$, where the parameter $\lambda$ can be viewed as a “coupling constant” and where it is understood that all field modes in the system are treated quantum mechanically. (Fig. 1 indicates the geometric layout, with a horizontally propagating system mode interacting with vertically propagating reservoir modes.)
The system mode is assumed to be in an arbitrary initial state described by a density matrix $\rho$ when it encounters the first beam splitter, at a moment when each element of the reservoir is in the same stationary initial state $\sigma$. In what follows, we establish under rather mild conditions on the common initial reservoir state $\sigma$, that the reduced system state $\rho_{K}$ that emerges from the $K$-th beam splitter asymptotically converges for large $K$ to a unique limiting state
@@
that generally depends on the coupling constant $\lambda$, but not on the initial state $\rho$ of the system mode. We then further establish that for weak coupling (i.e. small $\lambda$, corresponding to nearly perfect transmittance), the state $\rho_{\infty}^{\lambda}$ of the system mode that asymptotically emerges is precisely that “equipartitioned” state of thermal equilibrium having the same mean energy as each of the identically-prepared reservoir modes through which it has passed. We note that the dynamical process we consider is unitary and hence purely deterministic. Indeed, the only probabilistic element occurring in the model comes from that inherent in any quantum mechanical treatment. That is to say, no Stosszahlansatz is required for demonstrating this approach to equilibrium.
@@ 2 The Model
[…]
@@
We think of the family of modes $b_{k}$, each associated with an identical Fock space ${\cal H}_{k}={\cal H}_{b}$, as forming a reservoir and write ${\cal H}_{\mathcal{R}}={\cal H}_{1}\otimes\dots\otimes{\cal H}_{K}$ for the corresponding product Fock space. The reservoir is assumed to be initially in the $K$-fold product state $\sigma\otimes\dots\otimes\sigma$. The $a$-mode, that we think of as a small subsystem, has a corresponding Fock space ${\cal H}_{a}$. The small subsystem undergoes successive interactions with each of the modes $b_{k}$, and in what follows we establish properties of the asymptotic state
@@
$L(\rho)={\rm Tr}_{b}\big{(}S(\rho\otimes\sigma)S^{\dagger}\big{)},\qquad S={\rm e}^{\lambda(a^{\dagger}b-ab^{\dagger})},$ (6)
@@
$D_{b}(z)=\exp(zb^{\dagger}-z^{*}b),\qquad z\in\mathbb{C},$ (8)
@@
$A=2\begin{pmatrix}\text{Cov}_{\sigma}[b,b]&\text{Cov}_{\sigma}[b,b^{\dagger}]\\ \text{Cov}_{\sigma}[b^{\dagger},b]&\text{Cov}_{\sigma}[b^{\dagger},b^{\dagger}]\end{pmatrix},\quad\textrm{and}\quad\Delta=\begin{pmatrix}\langle b\rangle_{\sigma}\\ \langle b^{\dagger}\rangle_{\sigma}\end{pmatrix}.$ (12)
@@ 3.1 Convergence to and expressions for the asymptotic state $\rho_{\infty}^{\lambda}$
$\chi_{\infty}^{\lambda}(z):=\lim_{K\rightarrow\infty}\langle{L^{K}(\rho)D_{a}(z)}\rangle_{\rho}=\prod_{k=0}^{\infty}\langle{D_{b}(\sin\lambda\,[\cos\lambda]^{k}z)}\rangle_{\sigma},$ (16)
@@
$\langle a\rangle_{\infty}^{\lambda}\,:={\rm Tr}(a\rho_{\infty}^{\lambda})$ $=$ $-\partial_{z^{*}}\chi_{\infty}^{\lambda}(0)=\frac{\sin\lambda}{1-\cos\lambda}\langle b\rangle_{\sigma},$ (17)
$\langle a^{\dagger}\rangle_{\infty}^{\lambda}:={\rm Tr}(a^{\dagger}\rho_{\infty}^{\lambda})$ $=$ $\quad\partial_{z}\chi_{\infty}^{\lambda}(0)=\frac{\sin\lambda}{1-\cos\lambda}\langle b^{\dagger}\rangle_{\sigma}.$ (18)
@@
$-\partial_{z}\partial_{z^{*}}\chi_{\infty}^{\lambda}(0)=-\partial_{z}\partial_{z^{*}}\chi_{\sigma}(0)-\frac{\sin\lambda}{1-\cos\lambda}\partial_{z^{*}}\chi_{\sigma}(0)\frac{\sin\lambda}{1-\cos\lambda}\partial_{z}\chi_{\sigma}(0)+\partial_{z^{*}}\chi_{\sigma}(0)\partial_{z}\chi_{\sigma}(0).$
@@
${\rm Tr}(a^{\dagger}a\rho_{\infty}^{\lambda})+\frac{1}{2}=-\partial_{z}\partial_{z^{*}}\chi_{\infty}^{\lambda}(0),\quad{\rm Tr}(b^{\dagger}b\sigma)+\frac{1}{2}=-\partial_{z}\partial_{z^{*}}\chi_{\sigma}(0),$
@@
$=$ ${\rm Tr}\sigma(b^{\dagger}-\langle b^{\dagger}\rangle_{\sigma})(b-\langle b\rangle_{\sigma})+\frac{1}{2}=\mathrm{Cov}_{\sigma}[b^{\dagger},b].$
@@
$S^{\dagger}D_{a}(z)S=\exp\Big{[}(cz)a^{\dagger}-(cz)^{*}a)\Big{]}\exp\Big{[}(sz)b^{\dagger}-(sz)^{*}b)\Big{]}=D_{a}(cz)D_{b}(sz).$
@@ 3.2 Gaussian and non-Gaussian asymptotic system states and return to equilibrium
$=$ $\exp\big{[}\tfrac{1}{4}s^{2}\big{(}\sum_{k\geq 0}c^{2k}\big{)}Z^{T}\Omega^{T}A\Omega Z-s\big{(}\sum_{k\geq 0}c^{k}\big{)}\Delta^{T}\Omega Z\big{]}$
$=$ $\exp\big{[}\tfrac{1}{4}\,Z^{T}\Omega^{T}A\Omega Z-\tfrac{s}{1-c}\,\Delta^{T}\Omega Z\big{]}.$
@@
[…]
@@
Thus, this fully quantum mechanical model exhibits a “return to equilibrium”, in which the small system mode is driven to the same thermal equilibrium shared by the already equilibrated reservoir. Similar return to equilibrium processes are familiar from the open quantum systems literature, where one often considers the dynamics to be generated by a Hamiltonian $H=H_{S}+H_{R}+\lambda V$, consisting of a system term, a reservoir term, and an interaction term with coupling constant $\lambda$. Return to equilibrium for such systems, where the system has finitely many levels, the reservoir is thermodynamically large, and the coupling suitably small, was proven in [13, 14, 15, 16, 17] (see also references therein). The setup considered in the current manuscript is somewhat different. The dynamics of the $a$ mode (the system), given by Eq. (4), is that of a repeated scattering process, as the system is in contact with fresh reservoir elements sequentially in time (each element being a $b$ mode). The main idea of return to equilibrium remains nevertheless the same: a unique system mode being out of equilibrium constitutes only a small perturbation to the global equilibrium of the joint system-reservoir universe.
Our analysis above shows that we do indeed have return to equilibrium for all values of the coupling constant $\lambda$. We note that the dynamics of a class of closely related, so-called repeated interaction models, was analyzed in [18, 19, 20, 21], albeit for essentially finite dimensional system parts and only for weak coupling ($\lambda$ small).
@@
in which $p_{n}(x)=\sum_{j=0}^{n}{n\choose j}\frac{1}{j!}(-x)^{j}$ is the $n$th Laguerre polynomial [11], which is known to be the characteristic function of the state $|n\rangle$. It is straightforward to show that the state associated with (23) is then not Gaussian except when $n=0$. Indeed, expanding [^2] We have $\ln[\chi_{\infty}^{\lambda}(t)]=-\frac{1}{2}t^{2}+\sum_{k\geq 0}\ln[p_{n}(s^{2}c^{2k}t^{2})]$. By expanding $p_{n}(\epsilon)=1-n\epsilon+\frac{1}{2}n(n-1)\epsilon^{2}+O(\epsilon^{3})$ for small $\epsilon$, and then expanding the logarithm, one readily obtains the above mentioned expression. the logarithm $\ln[\chi_{\infty}^{\lambda}(t)]$ for small $t\in\mathbb{R}$, it is straightforward to establish that the Taylor coefficient of $t^{4}$ in that expansion is $-\frac{n}{2}s^{2}$, which implies that the asymptotic state $\rho_{\infty}^{\lambda}$ is not Gaussian for $n>0$.

### v2

@@ Abstract
Rigorous derivations of the approach of individual elements of large isolated systems to a state of thermal equilibrium, starting from arbitrary initial states, are exceedingly rare. This is particularly true for quantum mechanical systems. We demonstrate here how, through a mechanism of repeated scattering, an approach to equilibrium of this type actually occurs in a specific quantum system, one that can be viewed as a natural quantum analog of several previously studied classical models. In particular, we consider an optical mode passing through a reservoir composed of a large number of sequentially-encountered modes of the same frequency, each of which it interacts with through a beam splitter. We first analyze the dependence of the asymptotic state of this mode on the assumed stationary common initial state $\sigma$ of the reservoir modes and on the transmittance $\tau=\cos\lambda$ of the beam splitters. This analysis allow us to establish our main result, namely that at small $\lambda$ such a mode will, starting from an arbitrary initial system state $\rho$, approach a state of thermal equilibrium even when the reservoir modes are not themselves initially thermalized. We show in addition that, when the initial states are pure, the asymptotic state of the optical mode is maximally entangled with the reservoir and exhibits less nonclassicality than the state of the reservoir modes.
@@ 1 Introduction
There is continued interest (see for example [1, 2, 3, 4, 5, 6] and references therein) in the longstanding problem of how large systems, particularly quantum mechanical ones, undergo the ubiquitous process of thermalization, i.e., how it is that they are inevitably observed to approach a state of thermal equilibrium, starting from essentially arbitrary initial conditions. For large isolated quantum mechanical systems, much of this recent interest has focused on the difficult task of verifying the (weak or strong) Eigenstate Thermalization Hypothesis (ETH) of Deutsch in specific systems. According to the ETH, energy eigenstates of large systems tend, overwhelmingly, to have macroscopic properties consistent with the thermodynamically-equilibrated states that the systems are expected to approach. We refer to [1, 6] for overviews of this topic and its link with the problem of thermalization. In [6], the validity of the ETH in translationally invariant lattice systems is in particular discussed.
@@
In this paper we present a fully quantum mechanical model in which a similar mechanism of repeated scattering drives a single degree of freedom of a many-body system to thermal equilibrium. We show that the state of the single degree of freedom converges to an equilibrium state, even though the many-body environment it is coupled to, is itself not in equilibrium. We call this dynamical process “approach to equilibrium”. It is more intriguing than, and different from, the well-known process of “return to equilibrium”, whereby the single degree of freedom is driven to equilibrium when coupled to a many-body system which is itself in equilibrium.
The model we use belongs to a general class of models referred to as “collision models” or “repeated interaction models”. They have long been used as a versatile tool to efficiently model a variety of phenomena in equilibrium and non-equilibrium statistical mechanics, as well as in quantum information theory. An extensive overview of the use of and physical intuition behind such collision models can be found in the recent survey paper [12] and references therein. The mathematical formalism to analyse such models was developed in [13, 14, 15, 16].
In our model, a single optical field mode (the “system” mode) couples to a large reservoir of $K$ independent optical field modes of the same frequency through a sequence of $K$ identical beam splitters, each having a transmittance $\tau=\cos\lambda$, where the parameter $\lambda$ can be viewed as a “coupling constant” and where it is understood that all field modes in the system are treated quantum mechanically. (Fig. 1 indicates the geometric layout, with a horizontally propagating system mode interacting with vertically propagating reservoir modes.)
The system mode is assumed to be in an arbitrary initial state described by a density matrix $\rho$ when it encounters the first beam splitter, at a moment when each element of the reservoir is in the same stationary initial state $\sigma$. The reservoir is therefore initially in a product state. In what follows, we establish under rather mild conditions on the common initial reservoir mode state $\sigma$, that the reduced system state $\rho_{K}$ that emerges from the $K$-th beam splitter asymptotically converges for large $K$ to a unique limiting state
@@
that generally depends on the coupling constant $\lambda$, but not on the initial state $\rho$ of the system mode.
We then further establish that for weak coupling (i.e. small $\lambda$, corresponding to nearly perfect transmittance), the state $\rho_{\infty}^{\lambda}$ that the system mode asymptotically approaches is precisely that “equipartitioned” state of thermal equilibrium having the same mean energy as each of the identically-prepared reservoir modes through which it has passed.
We note that the dynamical process we consider is unitary and hence purely deterministic. Indeed, the only probabilistic element occurring in the model comes from that inherent in any quantum mechanical treatment. That is to say, no Stosszahlansatz is required for demonstrating this instance of approach to equilibrium. We further stress that, since the reservoir considered in this work is not in a thermal equilibrium state, the phenomenon highlighted here is not one of “return to equilibrium”, which is better understood and has been much more extensively studied, including for collision models [12, 17], and more generally through the use of Lindblad equations, in the weak coupling limit. There are also extensions to the non-equilibrium situation, where an open system is in contact with several reservoirs at different temperatures [18, 19]. The continuous time limit of a non-equilibrium collision model was analyzed in [20], for arbitrary coupling strength.
Return to equilibrium is intuitively understood as a stability result. The reservoir is supposed to be initially in equilibrium, while the probe degree of freedom, assumed to be weakly coupled to the reservoir, is not, so that the state of the full system can be viewed as a small perturbation of the global equilibrium state. At long times, the full coupled system then “returns” to equilibrium. Return to equilibrium occurs in the model we consider here as well. In fact, it is not limited to small coupling, but holds at all coupling strengths, as we will see below. Moreover, we do not resort to the continuous time limit of the repeated scattering model we consider. Rather, we analyze the discrete time collision model directly. To the best of our knowledge, the more elusive and difficult phenomenon of approach to equilibrium demonstrated here has not been previously shown to occur in collision models.
@@ 2 The Model
$S_{k}={\rm e}^{\lambda(a^{\dagger}b_{k}-ab_{k}^{\dagger})},\qquad\mbox{where\quad$0\leq\lambda\leq\tfrac{\pi}{2}$}$ (2)
@@
We think of the family of modes $b_{k}$, each associated with an identical Fock space ${\cal H}_{k}={\cal H}_{b}$, as forming a reservoir and write ${\cal H}_{\mathcal{R}}={\cal H}_{1}\otimes\dots\otimes{\cal H}_{K}$ for the corresponding product Fock space. The reservoir is assumed to be initially in the $K$-fold product state $\sigma\otimes\dots\otimes\sigma$. The $a$-mode, that we think of as a small subsystem, has a corresponding Fock space ${\cal H}_{a}$ and is initially in the state $\rho$. The initial state of the total system-reservoir complex is thus the product state $\rho\otimes\sigma\otimes\cdots\otimes\sigma$.
The small subsystem undergoes successive interactions with each of the modes $b_{k}$, and in what follows we establish properties of the asymptotic state
@@
$L(\rho)={\rm Tr}_{b}\big{(}S(\rho\otimes\sigma)S^{\dagger}\big{)},\quad S={\rm e}^{\lambda(a^{\dagger}b-ab^{\dagger})},$ (6)
@@
$D_{b}(z)=\exp(zb^{\dagger}-z^{*}b),\qquad z\in\mathbf{C},$ (8)
@@
$A$ $=$ $2\begin{pmatrix}\text{Cov}_{\sigma}[b,b]&\text{Cov}_{\sigma}[b,b^{\dagger}]\\ \text{Cov}_{\sigma}[b^{\dagger},b]&\text{Cov}_{\sigma}[b^{\dagger},b^{\dagger}]\end{pmatrix},\ \textrm{and}\quad$ (12)
$\Delta$ $=$ $\begin{pmatrix}\langle b\rangle_{\sigma}\\ \langle b^{\dagger}\rangle_{\sigma}\end{pmatrix}.$ (13)
@@ 3.1 Convergence to and expressions for the asymptotic state $\rho_{\infty}^{\lambda}$
$\chi_{\infty}^{\lambda}(z)$ ::=:= $\lim_{K\rightarrow\infty}\langle{L^{K}(\rho)D_{a}(z)}\rangle_{\rho}$ (17)
$=$ $\prod_{k=0}^{\infty}\langle{D_{b}(\sin\lambda\,[\cos\lambda]^{k}z)}\rangle_{\sigma},$
@@
$\langle a\rangle_{\infty}^{\lambda}\,$ ::=:= ${\rm Tr}(a\rho_{\infty}^{\lambda})$ (18)
$=$ $-\partial_{z^{*}}\chi_{\infty}^{\lambda}(0)=\frac{\sin\lambda}{1-\cos\lambda}\langle b\rangle_{\sigma},$
$\langle a^{\dagger}\rangle_{\infty}^{\lambda}$ ::=:= ${\rm Tr}(a^{\dagger}\rho_{\infty}^{\lambda})$ (19)
$=$ $\quad\partial_{z}\chi_{\infty}^{\lambda}(0)=\frac{\sin\lambda}{1-\cos\lambda}\langle b^{\dagger}\rangle_{\sigma}.$
@@
$-\partial_{z}\partial_{z^{*}}\chi_{\infty}^{\lambda}(0)=-\partial_{z}\partial_{z^{*}}\chi_{\sigma}(0)$
$-\frac{\sin\lambda}{1-\cos\lambda}\partial_{z^{*}}\chi_{\sigma}(0)\frac{\sin\lambda}{1-\cos\lambda}\partial_{z}\chi_{\sigma}(0)$
$+\partial_{z^{*}}\chi_{\sigma}(0)\partial_{z}\chi_{\sigma}(0).$
@@
${\rm Tr}(a^{\dagger}a\rho_{\infty}^{\lambda})+\frac{1}{2}$ $=$ $-\partial_{z}\partial_{z^{*}}\chi_{\infty}^{\lambda}(0),$
${\rm Tr}(b^{\dagger}b\sigma)+\frac{1}{2}$ $=$ $-\partial_{z}\partial_{z^{*}}\chi_{\sigma}(0),$
@@
$=$ ${\rm Tr}\sigma(b^{\dagger}-\langle b^{\dagger}\rangle_{\sigma})(b-\langle b\rangle_{\sigma})+\frac{1}{2}$
$=$ $\mathrm{Cov}_{\sigma}[b^{\dagger},b].$
@@
$S^{\dagger}D_{a}(z)S$ $=$ $\exp\Big{[}(cz)a^{\dagger}-(cz)^{*}a)\Big{]}$
×$\times\exp\Big{[}(sz)b^{\dagger}-(sz)^{*}b)\Big{]}$
$=$ $D_{a}(cz)D_{b}(sz).$
@@ 3.2 Gaussian and non-Gaussian asymptotic system states and return to equilibrium
$=$ $\exp\big{[}\tfrac{1}{4}s^{2}\big{(}\sum_{k\geq 0}c^{2k}\big{)}Z^{T}\Omega^{T}A\Omega Z$
$\qquad-s\big{(}\sum_{k\geq 0}c^{k}\big{)}\Delta^{T}\Omega Z\big{]}$
$=$ $\exp\big{[}\tfrac{1}{4}\,Z^{T}\Omega^{T}A\Omega Z-\tfrac{s}{1-c}\,\Delta^{T}\Omega Z\big{]}.\qquad$
@@
$A_{\infty}^{\lambda}=A\quad\mbox{and}\quad\Delta_{\infty}^{\lambda}=\frac{\sin\lambda}{1-\cos\lambda}\Delta.$
@@
Thus, this fully quantum mechanical model exhibits a “return to equilibrium”, in which the small system mode is driven to the same thermal equilibrium shared by the already equilibrated reservoir. Similar return to equilibrium processes are familiar from the open quantum systems literature, where one often considers the dynamics to be generated by a Hamiltonian $H=H_{S}+H_{R}+\lambda V$, consisting of a system term, a reservoir term, and an interaction term with coupling constant $\lambda$. Return to equilibrium for such systems, where the system has finitely many levels, the reservoir is thermodynamically large, and the coupling suitably small, was proven in [23, 24, 25, 26, 27] (see also references therein). The setup considered in the current manuscript is somewhat different. The dynamics of the $a$ mode (the system), given by Eq. (4), is that of a repeated scattering process, as the system is in contact with fresh reservoir elements sequentially in time (each element being a $b$ mode). The main idea of return to equilibrium remains nevertheless the same: a unique system mode being out of equilibrium constitutes only a small perturbation to the global equilibrium of the joint system-reservoir universe. Our analysis above shows that we do indeed have return to equilibrium for all values of the coupling constant $\lambda$.
@@
in which $p_{n}(x)=\sum_{j=0}^{n}{n\choose j}\frac{1}{j!}(-x)^{j}$ is the $n$th Laguerre polynomial [21], which is known to be the characteristic function of the state $|n\rangle$. It is straightforward to show that the state associated with (24) is then not Gaussian except when $n=0$. Indeed, expanding [^2] We have $\ln[\chi_{\infty}^{\lambda}(t)]=-\frac{1}{2}t^{2}+\sum_{k\geq 0}\ln[p_{n}(s^{2}c^{2k}t^{2})]$. By expanding $p_{n}(\epsilon)=1-n\epsilon+\frac{1}{2}n(n-1)\epsilon^{2}+O(\epsilon^{3})$ for small $\epsilon$, and then expanding the logarithm, one readily obtains the above mentioned expression. the logarithm $\ln[\chi_{\infty}^{\lambda}(t)]$ for small $t\in\mathbf{R}$, it is straightforward to establish that the Taylor coefficient of $t^{4}$ in that expansion is $-\frac{n}{2}s^{2}$, which implies that the asymptotic state $\rho_{\infty}^{\lambda}$ is not Gaussian for $n>0$.